OpenStudy (anonymous):

the perimeter of a triangle is 83inches. the longest side is three times the length of the shortest side and 17inches more than one-half the sum of the other two sides. use a system of equations to find the length of each side.

OpenStudy (anonymous):

a=3b a = 0.5(b+c) + 17 a + b + c = 83 so 3b = 0.5b + 0.5c + 17 and 3b + b + c = 83 2.5b = 0.5c + 17 4b = -c + 83 5b = c + 34 add 9b = 117 b = 13 so a = 3*13 = 39 and c = 83-39-13 = 31 lengths of sides = 13,31 and 39